OpenStudy (anonymous):

A train leaves Atlanta at 1:00PM. A second train leaves the same city in the same direction at 6:00 PM. The second train travels 50 mph faster than the first. If the second train overtakes the first at 10:00 PM, what is the speed of each of the two trains?

OpenStudy (anonymous):

9x=4(x+50) 5x=200 x=40 40.90

OpenStudy (a_clan):

v2 - v1 = 50 ...(1) v1*9 = v2*4 .....(2)

OpenStudy (anonymous):

train one is travelling at 40mph and train two is travelling at 90 mph

OpenStudy (anonymous):

so the first train is 41??? and the second would be 91??